OpenStudy (anonymous):

help for a medal Make a conjecture about the next item in the sequence. 6, 12, 9, 18, 15 a. 45 c. 12 b. 23 d. 30 @kx2bay @Jagking @Cjsmall14

OpenStudy (solomonzelman):

I see the pattern of times 2, minus 3, times 2, minus 3, and so on...

OpenStudy (anonymous):

It is a pattern. 6*2=12; 9*2=18 15*2= (blank). What is the last answer to the problem. 15*2=?

OpenStudy (solomonzelman):

\(\large\color{black}{ \bf 6 \times 2=12 }\) \(\large\color{black}{ \bf 12-3=9 }\) \(\large\color{black}{ \bf 9 \times 2=18 }\) \(\large\color{black}{ \bf 18 -3=15 }\) \(\large\color{black}{ \bf 15 \times 2... }\)